S-Lo (S Band Wide) Gregorian Dome Receiver Log

as of Friday, 03-Dec-2021 05:57:12 AST

Go to Most Recent Entry

ANS anolla@naic.edu 10/7/99 Jose Vives removed the horn of the antenna in order to check the new protection system.
ANS anolla@naic.edu 10/12/99 Receiver was shut off in order to warm it up. The tsys. has been degrading since July, so it is suspected of ice inside the dewar.
ANS anolla@naic.edu 10/14/99 Receiver was pumped for 48 hours , then cooled down . Amps were turned on but there is not an oportunity to check tsys. , they are using comp. terminals in the control room to run experiment in progress ( 430 Mhz xmmter world day )
ANS anolla@naic.edu 10/20/99 On 10/18/99 WG brought down the WS band control chassis to check out the amplifier LED drivers for the cryogenic amps , the front panel indicators were not working. He found that there was not drive voltage going to the dewar . He found inside the chassis that the +15 v was removed from the terminal strip , also the dewar LED return was removed , plus the +5 v regulator was dead.
ANS anolla@naic.edu 11/2/99 Channel A was tested with a 2270/100 filter in order to check if there was a relation with the tsys. increase and been without filters .The results were the same tsyst (54 deg) with or without filters in the sys.
ANS anolla@naic.edu 11/8/99 Receiver Dewar taken down to check its Tsys at the Receiver Test Range.
ANS anolla@naic.edu 1/27/2000 Receiver Dewar was installed on 1/25/2000 , and it was cooled down today.
ANS anolla@naic.edu 3/29/2000 New pneumatic autoshutter installed by Jose Vives to protect recvr from S band xmmter.
ANS anolla@naic.edu 4/17/2000 Autoshutter removed due to problems with it. Jose Vives found that air compressor was not working properly.
ANS anolla@naic.edu 5/17/2000 2380 Mhz (125 Mhz B.W.) filters were installed in the system to block R.F.I.
ANS anolla@naic.edu 5/18/2000 2380 Mhz filters removed from system.
ANS anolla@naic.edu 6/8/2000 The isolation with the shutter was checked and is 55 dbm.
ANS anolla@naic.edu 8/25/2000 The Tsys is deteriorating.From 55 k deg is up for 60 k deg , it will be taken down in order to check it at the RTR.
ANS anolla@naic.edu 9/12/2000 Receiver was taken down due that the Berkshire amps got burned out because the protective shutter get stucked not closing the feed during an S band xmtter run on 9/9/2000 .Amplifier model S-2.4-30H sn 102 was on chnn. A and sn 105 was on chnn.B. Both were shipping to Berkshire today in order to be fixed.
ANS anolla@naic.edu 10/5/2000 Receiver amps installed , back to Dome after amps installation.To be cooled down tomorrow.
ANS anolla@naic.edu 10/9/2000 Due to a power dip on Saturday afternoon compressor didn't started , warming up the receiver. Been cooled today.
ANS anolla@naic.edu 10/10/2000 Receiver cooled & checked for T Sys.Also the hybrid was installed for SETI people in order to polarized it.
ANS anolla@naic.edu 10/25/2000 W Genter removed the hybrid from receiver so it is lineal again.
ANS anolla@naic.edu 12/11/2000 Receiver taken down for maint. since 11/14/2000 , and today is at the receiver testing range building.
ANS anolla@naic.edu 12/28/2000 EC found at the RTR that chn B was unstable , have to open Dewar to check it.
ANS anolla@naic.edu 3/14/2001 Since mid Dec 2000 the receiver has been developing different problems at the RTR, chnn B is stable now, but chann A started an unstability .We found the input cables at the input of the amps at the Dewar broken, but didn't solve the unstability of chnn A. So we decided to replace the amp at chnn A . Amp sn 102 was there and was replaced with sn 114.
ANS anolla@naic.edu 3/16/2001 The chnn. A it's much better with amp sn 114, more stable. Still have some problems with the recovery at the high part of the band.
ANS anolla@naic.edu 4/3/2001 The Dewar crosshead removed due to contamination (water was found) , it is ED 024 (1020) , it will be replaced by ED 027.Still thinking of the displacer friction problem .
ANS anolla@naic.edu 4/3/2001 Amp sn 102 that was bad came back from manufacturer.
ANS anolla@naic.edu 4/4/2001 Amp at chann A replaced with amp sn 102, it was sn 114. Biases for Amp sn 102 are: Stage 1: 1.5V/5.2mA 2: 2.0V/6.0mA 3: 3.0V/10.0mA Biases for amp sn 114 are: Stage 1: 3.0V/5.0mA 2: 2.5V/7.0mA 3: 3.0V/10.0mA Both amps has a tsys about +/- 20k
ANS anolla@naic.edu 4/18/2001 The crosshead again contaminated, the cool down was started but after two hours got stocked.
ANS anolla@naic.edu 4/19/2001 Crosshead removed from Dewar, no visible contamination seen on it, except for some friction seen on both displacers. EC want to check timing on both 1020 crossheads available, ER to perform the timing check out. Meanwhile we are going to circulate the helium thru the system in order to remove if there is some.
ANS anolla@naic.edu 5/1/2001 The compressor helium was cleaned out thru the absolute filter in liquid nitrogen.
ANS anolla@naic.edu 5/10/2001 Today we try to install spare crosshead (ED 024) on system but something is wrong with it, it goes back and forth. Tried different crosshead driver and still bad. Also we have to modified the 16 stage seal (sanded a bit), it was too wide and didn't fit on displacer.
ANS anolla@naic.edu 5/16/2001 JR found that crosshead ED 027 valve driver was misaligned due to this the yoke assembly was rubbing against it and wearing out . The displacer 16 degree stage was out of round causing it to rub against the cylinder, this made the motor work harder. WR reduced the diameter of the both 16 degree displacers, they now go in smoothly. We installed the frig on the sband wide, it now works ok . we also diagnosed ED24 as having lost it's magnetism , ( due to bad displacer ??? ) it will be sent to Austin for repairs ($60)
ANS anolla@naic.edu 5/18/2001 The crosshead was getting stocked again, it was looking like contamination .So I called L Wray to notify it.Also I called EC to tell him about the problem , then he ask me if I played with the phase pot at the crosshead driver , I told him that yes but I tried again and almost at the lowest CCW position the crosshead started running normaly.Also I noticed grease at the driver ball bearing.
ANS anolla@naic.edu 5/22/2001 A microsw was installed at the horn to be a tell & tale of the ant shutter.
ANS anolla@naic.edu 5/23/2001 The system was checked by LW , WG and G.It is working properly.
ANS anolla@naic.edu 5/24/2001 Cal noise diodes are back , they are models Micronetics NS-213 . The Sn 7001 is installed in channel A and Sn 7002 in channel 2.
ANS anolla@naic.edu 5/31/2001 JR found that ND Sn 7001 was bad; it was generating noise with out drive. Have to be sending back to factory in order to be fixed.
ANS anolla@naic.edu 6/5/2001 Changed amplifier in Channel A. #102 was removed, #114 inserted. The bias levels have NOT been changed yet.
ANS anolla@naic.edu 6 june 2001 cooled down
ANS anolla@naic.edu 8 June 2001 crosshead noisy, bill genter turned off .
ANS anolla@naic.edu 11 June 2001 removed austing crosshead, greasy bearing , some binding of piston on cylinder. Coco and Cheo cleaned both lines up to compressor , pumped overnight.
ANS anolla@naic.edu 13 June 2001 Castro gave ok to install new CTI crosshead , cooled down ok , (less than 2 hours to close valve).
ANS anolla@naic.edu 14 June 2001Demo of rcvr to one telescope operation workshop group.
ANS anolla@naic.edu 19 June 2001Lisa, monro and cheo measured sky vs. eccosorb at 0700hrs. Chan A still ~ 10K higher than B.
ANS anolla@naic.edu 20 June 2001Lisa turned compressor off
ANS anolla@naic.edu 21 June 2001switched amplifiers: from Ch A:114 Ch B:105 to Ch A:105, Ch B:114
ANS anolla@naic.edu 22-23 June 2001cooled receiver.
ANS anolla@naic.edu 25 June 2001 noticed that the bias settings weren't set correctly for the previous week's measurements. ARG!
ANS anolla@naic.edu 8/27/2001 The system was installed back on 8/24/2001, almost after a year downstairs. The shutter was closed manually due to problems with its driver; the SB xmmter was on schedule.
ANS anolla@naic.edu 8/30/2001 Shutter is working properly except for the indicators, for some reason that we don't know the wiring was wrong. Indicators has to be rewired in order to indicate the shutter status. Also the WSB receiver ND chassis taken down in order to replace the bad ND.So the recvr rf is down, but cryo is up.
ANS anolla@naic.edu 10/19/2001 The post amps chassis was taken down for modifications on the rf system.
ANS anolla@naic.edu 10/24/2001 A 10 db D.C. was added in series with the post amp in order to have rf sampler of the rf signal from the Dewar , it will be of course 10 db down the main signal. The recvr was installed back and tested.
ANS anolla@naic.edu 11/5/2001 Cryo comp ED 007 replaced for service with comp ED 008.JR run the recvr with cryo comp ED 009 from 11/2/2001 until 11/5/2001 for testing the cryo comp, then replaced it with comp ED 008 on 11/5/2001.
ANS anolla@naic.edu 2/15/2002 For some reason that we don't know the cryocomp has been shut off a couple of times.
ANS anolla@naic.edu 2/25/2002 Cryocomp was shut off since yesterday , they turn it on but failed again .It is leaking oil. Have to replace it tomorrow.
ANS anolla@naic.edu 2/28/2002 The cryocomp ED 008 that was leaking oil was left, but it fail again and thanks God we pinpoint its problem with its cooling fan, it was getting stock. So we installed the spare cryocomp ED 010 (by the way we found that the on/off sw of these replacement was bad, it was not making contac, so we splice the wires of it).
ansanolla@naic.eduThursday, September 19, 2002 13:06:22Cryocomp 008 was replaced with cryocomp 010 (modified one).
ansanolla@naic.eduMonday, December 9, 2002 12:49:20Dewar turned off in order to replace cryopump.
ansanolla@naic edu.Wednesday, December 11, 2002 14:41:34Cryopump was replaced , but it failed to cool down the receiver.Was found getting stock , tried to tune it with cryopump driver but didn't work.
ansanolla@naic edu.Thursday, December 12, 2002 13:52:10Cryopump ED 028 that was installed on 12/11/2002 didnt work.So it was replaced again with cryopump ED 027.
ans anolla@naic edu.Friday, December 13, 2002 09:15:57Cryo temps are 16.4k & 88.6k .These is it after replacing cryopump for a second time.Have to check cryopump replaced.
aansanolla@naic edu.Thursday, December 19, 2002 15:59:26A filter bank was added to the system.
ansanolla@naic edu.Thursday, December 26, 2002 15:28:46The filter bank for the system was installed and checked.
ansanolla@naic.eduTuesday, April 8, 2003 12:51:38System hooked up for SETI people.Splitter added to its noise source thru the D.C.
ansanollaq@naic.eduMonday, May 12, 2003 14:51:32Due to cryosystem leak on the SBH , the cryosystem of the SBW was swap to it and viseversa.
ansanolla@naic.eduMonday, May 19, 2003 11:45:36The cryosystem still loosing 20 psi per day.Next maint we are going to switch cryocompin order to pinpoint leak.
ansanolla@naic.eduThursday, May 22, 2003 14:00:56Cryocompressor was replaced in order to pinpoint the 20 psi been lost daily due to the He leak . was cryocomp ED 008 now is ED 012.
ansanolla@naic.eduWednesday, May 28, 2003 09:46:26SBW cryogenic system shut off in order to fix a leak in the supply line at the S band xmmiter room.Cryocomp off , and the He lines are been monitored with meters to pinpoint the 20 psi/day leak.
ansanolla@naic.eduMonday, June 16, 2003 14:15:05The cryocomp ED 010 is back into the system.It was with the SBH system due to a He leak on it. It was at the flex cryo line in the top of the Dome, was the return line.
ansanolla@naic.eduMonday, April 12, 2004 15:01:44There is not shutter due to burned out air compressor.
ansanolla@naic.eduTuesday, April 13, 2004 15:54:21Yesterday during the rutine check out we found the air compressor for the shutters burned out. Today it was replaced with a new one. Shutters looks to be working properly.
ansanolla@naic.eduWednesday, April 28, 2004 16:09:22System was warming up ,was 19.9 at 16 and 125 at 70 . So Coco decide to hook up tha vacuum pump and it drop down to 15.9 & a 100 . I pumped it before but didn't work with me. Keeping an aye on it.
ansanolla@naic.eduMonday, May 10, 2004 09:02:00System down for modifications.
lswlwray@naic.eduFriday, May 14, 2004 10:55:28New calibration diodes: Noise/com diodes have replaced the old Micronetics ones.
ansanolla@naic.eduFriday, May 14, 2004 14:22:48System back to normal. Needs recalibration due to new noise cal diodes.
ansanolla@naic.eduFriday, May 14, 2004 14:25:13Focal point readjusted , is -2.6 " as it was.
ansanolla@naic.eduMonday, May 17, 2004 11:38:20Dewar cryo temp at 16 was 61. Also during a test to the cryopump , got locked , need replcement.
ansanolla@naic.eduMonday, May 24, 2004 09:23:10Cryopump was replaced (ED 027)for (ED 028), but it got locked again.On 5/21/2004 they removed the lock up cryopump. So receiver is down for a while.
ansanolla@naic.eduTuesday, May 25, 2004 15:42:55Coco replaced the cryopump ED 027 on 5/24/2004 by ED 028 , also he make the 16 deg displacer loose in order to make it swivel.But today we found that system was not working properly , for some reason the cryopump pop up valve blew lowering down the cryocomp preassure to 50 psi , why the cryocomp didn't turn off itself ,I dont know.But even with a new pop up valve still firing it .So Vives put a seal instead and that solve the problem.
lswlwray@naic.eduTuesday, June 1, 2004 16:06:38Monday May 31: When JR went upstairs to check the systems after the power outage, the cryopump was making bad sounds (mechanical misalignment or motor problems), so he turned it off. Receiver is warm.
lswlwray@naic.eduTuesday, June 15, 2004 08:30:07Dewar still warm. Using compressor as a spare to clean cryo lines.
lswlwray@naic.eduMonday, June 21, 2004 08:35:57Friday June 18: Hooked up the SBW to new compressor. Started cooldown around 11am. By 4:30 was down to 19K/90K and dropping. Slight mechanical sound, but hopefully it will last through the weekend.
lswlwray@naic.eduMonday, June 21, 2004 08:36:23Receiver stayed cold all weekend (!!!)
ansanolla@naic.eduWednesday, June 30, 2004 11:22:52System was calibrated yesterday.Also at the cryocomp the remote sw conn was replaced from a jumper wire to a connector.
ansanolla@naic.eduFriday, July 2, 2004 11:26:40System down due to power shut down (main power line shorted at the rotary joint).
ansanolla@naic.eduSaturday, July 3, 2004 16:32:13System been cooled , wit will be ready for tonight.
ansanolla@naic.eduWednesday, July 21, 2004 15:37:57For some reason the the interlock relay wasn't going , so we bypass it with it's bypass stick.
ansanolla@naic.eduFriday, July 23, 2004 16:14:32SB xmitter unable to run due that the shutter of the SBW was not closing.The problem was that the limit sw at the close side lost tension on it's lever.It was adjusted and worked fine.
lswlwray@naic.eduTuesday, July 27, 2004 13:40:36Receiver warmed up yesterday (July 26) when compressor ED012 failed. Receiver is being cooled today.
lswlwray@naic.eduTuesday, July 27, 2004 13:50:00Receiver warmed up on July 26 due to compressor ED012 failure (leaky return hose). Receiver put back onto ED016.
ansanolla@naic.eduTuesday, July 27, 2004 13:50:10System been cooled.
ansanolla@naic.eduWednesday, July 28, 2004 09:11:15System back to normal operation.
ansanolla@naic.eduWednesday, July 28, 2004 13:44:16New name for system SLow/SBW.
ansanolla@naic.eduWednesday, September 22, 2004 13:03:48JRV found the rf cable between recvr and rf cabinet broken , had broken connector at the recvr end.
ansanolla@naic.eduWednesday, September 22, 2004 15:06:36JRV did a quick fix on the channel A cable to the IFLO rack by reconnecting the shield back. This 13.5' long cable has ~10dB Insertion loss at 2.3GHz. needs to be replaced by Times or some other lower loss cable.
ansanolla@naic.eduMonday, October 11, 2004 15:55:22System was down due to power failure on Saturday , was cooled down by BG and Ganesh.
ansanolla@naic.eduTuesday, November 30, 2004 14:28:53The calibration system was not working properly in all the receivers at the Dome . It was due that the post amp chassis for the CBH and XB were off. These was loading the other recvrs calibrators.
ansanolla@naic.eduMonday, December 27, 2004 15:28:46System down since12/23/2004. The post amp chassis was removed due that the noise select board was causing the 60 Mhz rfi birdie that was getting into the other systems.
ansanolla@naic.eduTuesday, December 28, 2004 11:40:05System up again.The 60 Mhz birdie was killed on the noise injection board.U5D & U6D in pins grounded. , U7 A&B Vcc removed , and U1 & U2 inputs were terminated with a 10 K resistor to ground.
ansanolla@naic.eduWednesday, January 12, 2005 13:03:39Cryocomp ED 016 was swap with cryocomp ED 013.JVR make these in order to clean out the mess at the cryocomp area at the Dome.
ansanolla@naic.eduTuesday, February 22, 2005 16:25:31There it was a power shut down at the Dome due to a braker shut down at the substation that supplies the power to it.It affected several receivers because it last more than 30 minutes.The recvrs were 430 ,SBH,SBL and SBN.Vacuum pump on all of them.These happened on 2/18/2005
ansanolla@naic.eduFriday, May 27, 2005 13:16:59SBL receiver is down , its crosshead was getting stock.So it was warmed up in order to replace its crosshead.Off since yesterday afternoon.
ansanolla@naic.eduMonday, July 18, 2005 13:16:24System warmed up in order to take it down to replace crosshead displacer cilynder.
ansanolla@naic.eduTuesday, August 16, 2005 11:44:40Jose Vives found that the displacers holder was out for .050" from the vertical plane of the crosshead.Also the channel B 1 amplifier was open.So rf amp B was replaced with an spare one. It was sn 114 , now is sn 113.
ansanolla@naic.eduThursday, August 18, 2005 15:23:56System installed back after modification of cryohead displacer cylinder.Also Bershire rf amp B1 open.Was sn 114 ,now is 113.Cal cables replaced ,were bad.
ansanolla@naic.eduFriday, August 19, 2005 15:39:06Crosshead ED 028 got locked up at the displacer cylinder ,system was warming up. Warmed up to replace the crosshead , it is now ED 024.
ansanolla@naic.eduMonday, August 22, 2005 11:01:24Crosshead locked up again , suspect of contamination. It happened on 8/20/2005.Test Dewar to run the system for decontamination.
ansanolla@naic.eduFriday, August 26, 2005 13:04:29System cooled after third try with a crosshead.It locked up previus two.
ansanolla@naic.eduFriday, December 9, 2005 10:30:39System was calibrated by PP. The small adsorber used.
ansanolla@naic.eduWednesday, February 22, 2006 15:15:11The shutter plate was stock at the horn , removed from it,tested several times and worked ok.
ansanolla@naic.edu Thursday, July 6, 2006 12:50:05System was calibrated today by Felipe.
ansanolla@naic.eduWednesday, July 26, 2006 11:33:20Receiver been warmed up in order to replace its cryohead.
ansanolla@naic.eduFriday, July 28, 2006 11:20:13The system is being cooled down.We had problems with the cryom comps .First the comp ED 013 when started cooldown yesterday its power sw quit working ,then spare cryo comp ED 016 was used but after 3 hours of being running , it quit.Then the former comp of the system (ED 013) was fixed and was installed back.
ansanolla@naic.eduMonday, July 31, 2006 13:07:09System was cooled down , but this morning we found the crosshead getting stock.So we purged it and adjusted its phase.Keep an eye on it.
ansanolla@naic.eduTuesday, August 1, 2006 11:17:07Cryohead is working fine after tuning it with the driver phase adjusment pot.
ansanolla@naic.eduTuesday, August 1, 2006 11:19:07Forgot to mention that cryo head ED 028 replaced the ED 029.
ansanolla@naic.eduWednesday, August 2, 2006 14:55:14System cryohead getting stock,CR working on it.
ansanollaMonday, November 6, 2006 10:42:55Cryohead got stock.CR & VN working on its replacement.
ansanolla@naic.eduTuesday, November 7, 2006 11:38:57CR working on the cryohead that was on system (ED 028), in order to install it back.
ansanolla@naic.eduThursday, November 9, 2006 13:01:14System being cooled down.CR installed cryo head ED 034 on it.Cryo head ED 028 down for repairs.
ansanolla@naic.eduFriday, November 10, 2006 14:11:11System back to normal operation.Cryo head is working properly as far as we know.
ansanolla@naic.eduThursday, December 28, 2006 13:27:52There it was a problem with these receiver during last week-end.Receiver was dead , Ganesh came to work with the ALFA rotation system last Sunday and also he found that the power sw of the bias chassis of the SBL was off,also the chann B diode exiter sw.Problaby when we were taken the recvrs readings last Friday accidentally we turn off the power sw of the bias chassis. We always ask for a Tsys all to double check that all the receivers are working properly , but that day the rotary floor was not available , so we told the telescope operator to perform the Tsys later.It was done at AST 14:06 hours. And it showed that the SBL was dead.But they didn't notified us.
ansanolla@naic.eduMonday, April 23, 2007 13:18:47System turned off due to painting project.
ansanolla@naic.eduWednesday, April 25, 2007 13:03:20System cryo comp taken down for service.
ansanolla@naicMonday, May 7, 2007 10:34:26CR replaced the bias box power , heather & rf amp power sw's with the locking type to avoid accidental throw of them. Also he is working in the new driver board for the rf amps exiter LED's.
crvWednesday, January 30, 2008 14:26:10The system was vac. and is cooling down
crvWednesday, February 6, 2008 14:03:51Reciver was swaped with recievers sbn and xb. The system was originalywith compressor ED009 and is now with compressor 008.
Dana Whitlowdwhitlow@naic.eduThursday, February 14, 2008 14:36:23Swapped crosshead driver with that of SBN, on 13 Feb 2008. SBN crosshead was in distress, making lots of humming noises and sometimes chattering and freezing.
crvThursday, February 28, 2008 14:28:26The reciever had problems with some calibrations on channel A. We replaced a rf swicht in channel A. The Tsys values for Ch a are normal now.
crvThursday, March 20, 2008 16:23:21The comp. Ed 009 was replaced with comp. Ed 013 Receiver is cold.
crvWednesday, April 30, 2008 10:53:46The shutter was not closing completely. We thigthen a loose screw and replaced a missing screw. Shutter working fine.
crvMonday, May 12, 2008 10:50:30The deawr is being vacuumed.for this weeken experiment.
crv Wednesday, June 18, 2008 16:02:42the dewar has started cooldown .
crvMonday, July 21, 2008 11:03:55system warmed up and compressor used on CBH.
Dana Whitlowdwhitlow@naic.eduFriday, October 10, 2008 08:50:15New CAL values were installed 29 Sept 08.
CRVMonday, October 13, 2008 14:39:32ON October the 10th 2008 Comp 9 was replaceed by comp.15 because it was goin to be used on CBH.
Stephen Jensensjensen@naic.eduWednesday, February 18, 2009 12:54:17Had to completely warm up SBW due to heavy ratcheting. Cleaned the pistons and cylendar & found moisture on the cylendar.
crvcrvMonday, August 10, 2009 10:17:37dewar started vac. process today at 9:30 am . The system was warmed due tothe power failures of wednesday august 5th 2009.
crvFriday, August 14, 2009 16:09:28Dewar cool down started today on comp.14 that had the fan replaced .
crvThursday, August 20, 2009 14:54:41system was warmed down on monday august 17 2009 due to racheting.
crvThursday, August 20, 2009 14:56:04cryo pump was removed and cilinder cleaned moister removed and started cooling down onwednesday august19 2009
crvThursday, August 20, 2009 14:56:42system is cold and running fine. Vac. pump removed.
crvcrios@naic.eduThursday, March 4, 2010 11:18:39Dewar Cryo pump was turned off because the temperature was raising and the VPG value was 35. When the Cryo pump sight glass was checked evidence of Bronze dust of the bushing were covering the glass due to bushing wear cryo pump running since January 2008. Cryo pump will be replaced once the Dewar temp. reach room temp.
crioscrios@naic.eduWednesday, March 10, 2010 15:07:08Receiver has started cooldown today the new cryopump is Ed27.
crvcrios@naic.eduTuesday, March 16, 2010 14:59:46system is warming up again due to contamination resudue of the last pureged made to the system yesterday.
crvcrios@naic.eduMonday, March 22, 2010 14:18:06The cryo pump was remove and replaced for a new one and started cooldown at 10:00 am. The removed cryo pump had a defective motor.
Denisdurbain@naic.eduFriday, October 5, 2012 11:28:56The shutter developped a small leake on one of the pressurized tube. The tube was contacting the solenoid valve and the heat from the motor when the shutter is in the closed position melted the plastic.Repair kit has been ordered
Denis Urbaindurbain@naic.eduThursday, October 11, 2012 08:13:53The shutter was repaired, new tube and stainless steel fittings were installed.
Denis Urbaindurbain@naic.eduMonday, February 25, 2013 13:45:19The compressor that was connected to the receiver was swapped with the compressor from the S band narrow. After the swap the coldhead started to make ratcheting noise. The coldhead was turned off and the lines disconnected. After the warm-up the coldhead was removed and water was found on the displacer that was also quite dirty. The Coldhead was refurbished by Guarionex while the vacuum was restored. After the installation the receiver was cooled, the second stage got cold very quickly and the vacuum pump could be disconnected early
Denis Urbaindurbain@naic.eduThursday, June 6, 2013 08:36:15The receiver had to be warmed up because of compressor problems. The C Band Hi was required to complete an astronomical program therefor the S Lo refregerator was stopped. At the moment we are running the 600 compressor at ful capacity, the S Radar, C Hi and S Lo are all connected to the same compressor. The S Lo 2nd stage temperature is around 20K and the Tsys 4-5K higher than normal, the C Hi Tsys is only 1-2K higher than normal. We will try to keep both receivers cold as long as the astromers accept the increase Tsys.
Felix O. Fernandezfelix.fernandez@naic.eduThursday, December 26, 2013 15:02:52The SBW/L Rx was recording temperatures of 120K at the OMT and the 20K stage temperature seemed unstable. It was decided to warm up the Rx on DEC-23-2013. The cold-head was removed and serviced on DEC-24-2013. Finally, the cold-head was replaced and the Rx started the cool-down process on DEC-26-1013. NOTE: The SBW/L RX dewar's vacuum sensor appears to be off by around 15 ATM. This might need replacement whenever the dewar is taken down for service, for now the vacuum sensor on the vacuum pumps me be used as a substitute.
Carlos Rioscrios@naic.eduThursday, March 20, 2014 15:18:00System cryo pump was replaced and colling started.
Cold-head contaminationfelix.fernandez@naic.eduTuesday, April 22, 2014 17:19:53Today,since the compressor driving the SBN, SBL and CBH was leaking He, it was decided to move the SBN and SBL receivers to the M600 compressor that typically is used for ALFA (note: ALFA is currently down for repairs). After 15-20 minutes of operation the SBN and SBL temperatures started rising. Once we got to the G. Dome we found both cold-heads to be severely contaminated. Thus, a thermal cycle was started today afternoon at 15:30.
Compressor failurefelix.fernandez@naic.eduWednesday, July 30, 2014 11:09:15The SBN, SBL/W , and CBH compressor (model M600) went "dead" after last Jun-19-2014 power failure. After several attempts to fix the problem while the compressor was upstairs it was concluded that the compressor's scroll pump was seized and was tripping the over-current sensor. Hence, SBN, SBL/W and CBH have been warm since that event. Today we finally got the opportunity to replace the compressor. All the aforementioned receivers will begin their vacuum and cool-down cycles sequentially today.
Cool-down cycle startedfelix.fernandez@naic.eduTuesday, August 5, 2014 13:12:11Today we started the cool-down cycle for the SBN cold-head. The vacuum gauge read 6.5 mTorr at the time we started the cold-head.
Cryo pump replacementcrios@naic.eduTuesday, April 14, 2015 15:29:30System starting to warmup for a cryo pump replacement. System cooling is poor.
Pol-B LNA Stg1 No Idfelix.fernandez@naic.eduThursday, October 13, 2016 10:29:52On 12-OCT-2016 at around 00:00 Pol-B's LNA lost current to its first stage. After checking the bias card and dewar-to-bias chassis cable we concluded that the problems lies inside the dewar. Hence, SBW must come down for repairs.
Thermal Cyclefelix.fernandez@naic.eduWednesday, January 16, 2019 10:22:55During 12-jan-19 there was a power dip and an unexpected thermal cycle started on SBN,SBW,LBW. Seems to be compressor related.
PolA LNA replacementfelix.fernandez@naic.eduMonday, January 28, 2019 10:37:30SBW's Pol-A LNA had a bad 1st stage. The gate became unresponsive and Id was no longer controlled by the bias circuit. On 201-01-28 the LNA (serial 105) was replaced by its spare (serial 102). Both LNAs are Berkshire Technologies Inc. Model S-2.4-30H. During the LNA replacement we also repaired the 20K temperature diode which had a faulty connection.

Back to Top